### v2

# sample.bib

Abstract
Purpose: We present an updated study evaluating the performance of large language models (LLMs) in answering radiation oncology physics questions, focusing on the recently released models.
Methods: A set of 100 multiple choice radiation oncology physics questions, previously created by a well-experienced physicist, was used for this study. The answer options of the questions were randomly shuffled to create ”new” exam sets. Five LLMs (OpenAI o1-preview, GPT-4o, LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet) with the versions released before September 30, 2024, were queried using these new exam sets. To evaluate their deductive reasoning capabilities, the correct answers in the questions were replaced with ”None of the above.” Then, the explaining-first and step-by-step instruction prompts were used to test if this strategy improved their reasoning capabilities. The performance of the LLMs was compared with the answers from medical physicists.
Results: All models demonstrated expert-level performance on these questions, with o1-preview even surpassing medical physicists with a majority vote. When replacing the correct answers with ’None of the above’, all models exhibited a considerable decline in performance, suggesting room for improvement. The explaining-first and step-by-step instruction prompts helped enhance the reasoning capabilities of the LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet models.
Conclusion: These recently released LLMs demonstrated expert-level performance in answering radiation oncology physics questions, exhibiting great potential to assist in radiation oncology physics training and education.
1 Introduction
Large language models (LLMs) have advanced rapidly. On the one hand, the size of the data used for pre-training and the number of model parameters have grown significantly [gururangan-etal-2020-dont, kaplan2020scalinglawsneurallanguage]; on the other hand, the fine-tuning methods and prompt engineering strategies have advanced substantially [ouyang2022traininglanguagemodelsfollow, hu2021loralowrankadaptationlarge]. Furthermore, agents and Retrieval-Augmented Generation (RAG) systems built on LLMs have seen considerable progress [xi2023risepotentiallargelanguage, lewis2021retrievalaugmentedgenerationknowledgeintensivenlp].The notable recent developments include OpenAI o1-preview [o1_webpage], GPT-4o [openai2024gpt4ocard], LLaMA 3.1 (405) [grattafiori2024llama3herdmodels], Gemini 1.5 Pro [geminiteam2024gemini15unlockingmultimodal], and Claude 3.5 Sonnet [claude_webpage], each demonstrating remarkable improvements in language comprehension, reasoning, and domain-specific expertise.
The rapid evolution of LLMs also renders prior performance evaluations outdated. As some LLMs cease providing services, new models are introduced, and existing versions are updated, studies published before may no longer accurately reflect the current state of LLM capabilities. A fresh evaluation is needed to address the dynamic landscape due to the rapid LLM development.
Although,large language models (LLMs) have shown great potential for applications in healthcare recently [liuzhengliang2023, Liu2024a, LIU2023100045, Liu2022SurveyON, Rezayi2022, zhangkai2024, zhang2024, Dai2023ChatAugLC, HOLMES2024e515, XIAO2024102204, Wu2023ExploringTT, LIAO2024128576, li2024, dai2023samaug], their direct use in specialized fields such as radiation oncology presents challenges and necessitates thorough evaluation and validation. Radiation oncology is distinct from other clinical practices due to two salient features: a high degree of complexity and stringent precision requirements [dawson2007advances, xing2006overview]. This is particularly true for advanced radiation therapy modalities like intensity-modulated X-ray-based radiation therapy and intensity-modulated proton therapy [quan2013preliminary, cao2012uncertainty, zaghian2017comparison, liu2018impact, an2017robust, liu2020robust, shan2020intensity]. Therefore, assessing the performance of LLMs in addressing questions related to radiation oncology physics is crucial for their successful integration into this field. Such evaluations not only show us how effectively they process and understand radiation oncology physics, but also help us understand their limitations. This provides us with guidance in this field, especially for their potential utilization in training and education in radiation oncology physics. Built on our prior work [10.3389/fonc.2023.1219326], we provided an updated study with refined methods, offering insights into the performance of the recently released LLMs in radiation oncology physics.
We utilized the 100-question radiation oncology physics exam we developed based on the American Board of Radiology exam style [liu2024radiationoncologynlpdatabase], and randomly shuffled the answer options to create ”new” exam sets. We then queried the LLMs with these new exam sets and checked their capability to answer questions accurately. We also evaluated their deductive reasoning capabilities and tested whether the explaining-first and step-by-step instruction prompts would improve their performance in answering radiation oncology physics questions.
2 Methods
The 100 multiple-choice exam questions on radiation oncology physics were created by our experienced medical physicist, following the official study guide of American Board of Radiology. That exam includes 12 questions on basic physics, 10 questions on radiation measurements, 20 questions on treatment planning, 17 questions on imaging modalities and applications in radiotherapy, 13 questions on brachytherapy, 16 questions on advanced treatment planning and special procedures, and 12 questions on safety, quality assurance (QA), and radiation protection. 17 out of the 100 questions are math-based and require numerical calculation.
All the evaluated LLMs were queried with the exam questions through Application Programming Interface (API) services provided by their respective hosts, except LLaMA 3.1 (405B), an open-source LLM, which was hosted by us locally at our institution. All the LLMs used were the recently released version before September 30, 2024. The temperature was set to 0.1 for all LLMs to minimize variability in their responses [^1] The temperature was set to 0.1 rather than 0 due to the different meanings of a temperature of 0 across different LLMs., with the exception of the OpenAI o1-preview, whose temperature was fixed at 1 and could not be changed by the user.
2.1 Randomly shuffling the answer options
A Python code was developed to randomly shuffle the answer options for the 100 multiple-choice questions five times. For each shuffle, we obtained a ”new” 100-question multiple-choice exam set. We then queried all the LLMs five times. For each time (Trial 1 - Trial 5), we queried all the LLMs with a ”new” exam set generated by randomly shuffling. Each question of every exam set was queried individually. We checked the distribution of the correct answers’ locations for the five exam sets and confirmed that the distribution of the correct options is fairly random among A, B, C, D, or E (only 2 questions offered option E).
The prompt we used for all the queries was as follows:
”Please solve this radiation oncology physics problem:
[radiation oncology physics problem]. ”
This allowed the LLMs to reason and answer freely. Table 1 illustrates an example of the trials and how we queried the LLMs with the questions. For the responses generated by the LLMs, we utilized the LLaMA 3.1 (405B) model hosted locally to further extract the chosen answer options (letter A, B, C, D, or E) from responses. We then conducted manual verification and correction of the extracted options and obtained the final answer option sheet for all LLMs. The accuracy of each LLM was reported as the mean score across the five trials, and the measurement uncertainty was reported as the standard deviation of the five trials, as shown in the following equations:
$\text{Mean}(\bar{x})=\frac{1}{N}\sum_{i=1}^{N}x_{i},$
$\sigma=\sqrt{\frac{1}{N-1}\sum_{i=1}^{N}(x_{i}-\bar{x})^{2}},$
where $x_{i}$ represents each measurement.
The results of the LLMs’ test scores were compared with the majority vote results from a group of medical physicists conducted in our previous study. The medical physicist group consisted of four experienced board-certified medical physicists, three medical physics residents, and two medical physics research fellows. For each question, the most common answer choice was selected as the group’s answer. In case of a tie, one of the most common answer choices was chosen randomly.
Table 1: Illustration of the prompts and questions of randomly shuffled options to evaluate LLMs’ performance on answering radiation oncology physics questions.
Trail Prompt Question
Trail 1 Please solve this radiation oncology physics problem:
Which of the following particles cannot be accelerated by an electric field?
(a) Neutrons
(b) Protons
(c) Electrons
(d) Positrons
Trail 2
Which of the following particles cannot be accelerated by an electric field?
(a) Protons
(b) Neutrons
(c) Electrons
(d) Positrons
Trail 3
Which of the following particles cannot be accelerated by an electric field?
(a) Positrons
(b) Protons
(c) Electrons
(d) Neutrons
Trail 4
Which of the following particles cannot be accelerated by an electric field?
(a) Electrons
(b) Neutrons
(c) Protons
(d) Positrons
Trail 5
Which of the following particles cannot be accelerated by an electric field?
(a) Electrons
(b) Positrons
(c) Neutrons
(d) Protons
2.2 Evaluating deductive reasoning capability
Deductive reasoning capability refers to the cognitive process of logically analyzing information to draw specific conclusions from general premises. The multiple-choice question with the answer option ”None of the above” can effectively evaluate the test-taker’s deductive reasoning capability, as it involves evaluating each option based on the information provided and ruling out incorrect choices contradicting known facts or logical outcomes to reach the correct answer. We therefore had replaced the correct option in the exam with ”None of the above.” Since transformer-based LLMs predict the next word based on prior contexts, changing the correct option to ”None of the above” removes a straightforward cue that might guide the model toward a known or patterned solution, thus forcing the LLMs to rely more on reasoning about the specific question and its options, rather than using surface-level lexical or statistical patterns that it may have learned.
2.2.1 Replacing the correct option with ”None of the above”
A Python code was developed to replace the correct option with ”None of the above” for all questions in the five new exam sets derived by random shuffling. For each trial (Trial 1 - Trial 5), we queried the LLMs with a set of exams in which both the correct option was ”None of the above,” and its location was randomly shuffled. We used the same prompt as in Sec. 2.1 across all queries, and each question in an exam set was queried individually. This setup challenges the LLMs to avoid pattern-based answering and not rely on any single choice, but to process each answer option by reasoning.
Similar to before, we utilized the previously described processes for answer-option extraction, manual verification, and correction, as outlined in Sec. 2.1. The performance accuracy and uncertainty of each LLM were reported as the average score and standard deviation across all five trials. Due to this setup, these exams were not used to test humans, as this pattern can be easily recognized by human test-takers [^2] Since each question was queried through the API individually, it is assumed that LLMs would not notice this pattern.
2.2.2 Explaining first and step-by-step instruction
To further check if explicitly asking the LLMs to explain first and then develop answers step-by-step would improve their deductive reasoning capability [wei2023chainofthoughtpromptingelicitsreasoning], we engineered the following prompt and queried the LLMs again with it:
”Please solve this radiation oncology physics problem:
[radiation oncology physics problem]
Please first explain your reasoning, then solve the problem step by step, and lastly provide the correct answer (letter choice).”
We used the five exam sets and conducted the querying process both as described in Sec. 2.2.1. All five LLMs were evaluated using this prompting strategy. Accuracy and uncertainty were reported as before. The results from this strategy were compared with the test results from original prompts, where no explanation or step-by-step answering was required, as described in Sec. 2.2.1.
3 Results
3.1 Results of exam sets with randomly shuffled options
The evaluation results of the exam sets with randomly shuffled options are presented in Fig. 1, where the height of each bar represents the mean test score, and the error bars indicate the standard deviations across five trials. All five LLMs exhibited strong performance, achieving mean test scores above 80%, which suggests their performance on these exams is comparable to that of human experts. When compared to the majority vote results from the medical physics group, the OpenAI o1-preview model outperformed the medical physicists with a majority vote. For math-based questions, both the o1-preview and GPT-4o models surpassed the medical physicists with a majority vote.
Figure 1: Evaluation results using five exam sets, where the answer options of the questions were randomly shuffled. Figure (a) illustrates the evaluation results for all questions and math-based questions, while Figure (b) presents the evaluation results broken down by different topics.
The raw counts of incorrect answers by the LLMs are shown in Fig. 2, where each color represents the incorrect answers by an LLM across trials. As observed, each LLM exhibited variability in answering questions across trials. Notably, the models also showed similarities in incorrectly answering certain questions. We analyzed the questions that all five models had ever answered incorrectly across all five trials, i.e., question numbers: 14, 27, 33, 42, 67, 87, 95, 96. Interestingly, only one of these questions was math-based, while the remaining seven were closely related to clinical medical physics knowledge, such as American Association of Physicists in Medicine (AAPM) Task Group (TG) reports and clinical experience. This observation suggests that current LLMs may still struggle with answering clinically-focused radiation oncology physics questions. For example, question number 42 does not involve any calculations but instead focuses primarily on clinical hands-on experience.
Figure 2: Distribution of incorrect answer counts of every LLM across five exam sets, where answer options were randomly shuffled. Questions that were correctly answered by LLMs in all five trials are not displayed in this figure.
3.2 Results of LLMs’ deductive reasoning capability
Fig. 3 shows the results of the deductive reasoning capability tests, where the correct answer options were replaced with ”None of the above” in all questions. Overall, all LLMs performed much more poorly compared to the results in Sec. 3.1. Given that transformer-based LLMs [vaswani2023attentionneed] were designed to predict the next word in a sequence, replacing the correct answers with ”None of the above” would likely disrupt their pattern recognition capabilities, thereby reducing their overall scores performed on the exam sets. Nonetheless, the OpenAI o1-preview and GPT-4o still outperformed the others, especially on math-based questions, indicating strong reasoning capabilities of these two models.
Figure 3: Deductive reasoning capability evaluation results for every LLM. The correct answer options were replaced with ”None of the above” in all questions. Figure (a) illustrates the evaluation results for all questions and math-based questions, while Figure (b) presents the evaluation results broken down by different topics.
Fig. 4 compares the performance of LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet models with the original simple prompts and with the explaining-first, step-by-step instruction prompts. Overall, all three models demonstrated improved reasoning capabilities with the latter prompting strategy. Notably, Gemini 1.5 Pro showed significant gains on math-based questions, increasing its score from 24% to 68%. The o1-preview and GPT-4o improved by only about 1% overall, too small to be represented in this figure.
Figure 4: Improvements in reasoning capabilities observed using the explaining-first and step-by-step instruction prompts for the LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet. Figure (a) illustrates the improvements for all questions and math-based questions, while Figure (b) presents the improvements broken down by different topics. (Note: The o1-preview and GPT-4o showed only about 1% overall improvement, too small to be represented in this figure.)
4 Discussion
4.1 Improvement of performance on answering radiation oncology physics questions of the state-of-art LLMs over the past two years
Over the past two years, our studies have observed a notable improvement in the performance of state-of-the-art LLMs on this highly specialized task - answering radiation oncology physics questions, as shown in Fig. 5. Earlier versions of ChatGPT, like GPT-3.5 in late 2022 [chatgpt_webpage], scored around 54%, showing clear gaps in domain-specific knowledge. With the introduction of GPT-4 in early 2023 [openai2024gpt4technicalreport], performance jumped to around 76%, reflecting improvements in accuracy and understanding. Subsequent releases of the GPT-4o model and more recently the o1-preview (both in 2024), pushed scores even higher to 90% and 94% respectively, indicating improved domain knowledge in radiation oncology physics. This steady improvement can be attributed to more extensive domain pre-training, refined architectural updates [brown2020languagemodelsfewshotlearners], and enhanced fine-tuning techniques, all of which have led to stronger reasoning skills, improved understanding, and better alignment with expert-level knowledge. The evolution of these models over the last two years underscores the rapid growth of LLMs, suggesting their potential as useful tools in areas such as radiation oncology physics training and education.
Figure 5: Growth of the state-of-art LLMs’ performance in radiation oncology physics over the past two years. Two dotted lines mark the average score and the majority vote score of the medical physicists.
4.2 Utilization of LLMs in radiation oncology physics training and education
With the excellent performance of LLMs in answering radiation oncology physics questions evaluated in this study, it offers a promising avenue for their utilization in radiation oncology physics training and education to enhance learning efficiency and accessibility. For example, LLMs can assist in educating graduate students, postdocs, and medical physics residents as they prepare for challenging medical physics certification exams, such as the American Board of Radiology certification exams. LLMs have the unique advantage of providing immediate and context-specific explanations of medical physics concepts and are efficient at retrieving information from a wide range of topics. In busy clinical environments, where mentors may sometimes be occupied with clinical duties, LLMs can serve as a useful tool to address easy questions and have learners pay more attention to more complex problems. This leads to more focused and productive discussions with mentors. In summary, LLMs can bridge gaps in traditional learning settings and support self-directed study in radiation oncology physics.
5 Conclusion
We evaluated recently released LLMs using a method that randomly shuffled the answer options of radiation oncology physics questions. Our results demonstrated that these models achieved expert-level performance on these questions, with some even surpassing human experts with a majority vote. However, when the correct answer options were replaced with ”None of the above,” all models exhibited a steep decline in performance, suggesting room for improvement. Employing the technique of explaining-first and step-by-step instruction enhanced the reasoning capabilities of LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet. The performance of these recently released LLMs exhibited great potential for assisting in radiation oncology physics training and education.
6 Acknowledgments
This research was supported by the National Cancer Institute (NCI) R01CA280134, the Eric & Wendy Schmidt Fund for AI Research & Innovation, The Fred C. and Katherine B. Anderson Foundation, and the Kemper Marley Foundation.
[…]